# DiVerify: Hardening Identity-Based Software Signing with Programmable Diverse-Context Scopes

### v2

@@ 5.3 Integrating Scope Providers
DiVerify requires scope providers must be able to authenticate themselves to the daemon and prevent scope replay to ensure that claims are both trustworthy and fresh. however, not all providers natively support these. To accommodate this, we introduce two strategies: wrapping and TEE integration—which support various provider capabilities. We discuss different categories of these providers and how they are supported in DiVerify.
Verifiable Providers (VP)
@@
are able to cryptographically authenticate themselves and issue scopes that are bound to a specific context, making them verifiable and resistant to replay. For example, OIDC tokens signed by a known Identity Provider such as Google. VPs are natively supported in DiVerify with minimal changes, and transferring their scopes only requires defining a syntax for policy verification.
Static Providers (SP)
@@
SPs do not support nonce injection, but the authenticity of the provider can still be verified through cryptographic signatures or secure communication channels. SPs are often stateless protocols that provide a fixed value whenever called. For example, this is the case of hardware keys using device attestation for PIV, where the same x509 certificate is provided every time an attestation request is made. This introduces two risks, first, without state-keeping an attacker can cache scope proofs. Second, if the value is consistent, then an attacker could also re-generate that value in another context.
We devise two strategies to support SPs:
@@
  • Zk-based SP: the scope provider is wrapped inside a process that provides a proof of the value being a correct device fingerprint. This can be transparently verified by both the verifier and the diverify daemon, and the nonce utilized can match the nonces provided to other scope providers.
  • nonce-hash SP: in some cases, authentication is enough and its only necessary to prove the scope requested was part of a signing event. In this case the reply is hashed along an input nonce and a triple (noncenonce, measurementmeasurement, sigsig(hashhash(noncenonce|measurementmeasurement)) is returned.
Opaque Providers (OP)
@@
OPs allow nonce injection to prevent replay attacks, but lack a reliable mechanism to verify the provider’s identity making them susceptible to spoofing if used in isolation. For example, this is the case of TPM ME and device fingerprinting writ large.
@@
Untrusted Providers (UT)
UTs support neither nonce injection nor provider verification, offering no protection against replay or impersonation attacks and representing the weakest security posture. We combine the strategies for OPs and SPs to support UTs.
@@
5.4 Distribution of DiVerify Policy
DiVerify’s security relies on the tamperproof distribution of its policies, which serve as the root of trust for verification. To ensure freshness, authenticity, and integrity, we use The Update Framework (TUF) [43].
TUF uses separation of duties to isolate signing different pieces of metadata about DiVerify policies, as illustrated in Figure˜3. The idea is that while machines can be trusted to sign with online keys—or keys sitting on the repository—higher-level statements about whether there are new or updated DiVerify policies (i.e., the timestamp and snapshot metadata), only humans are allowed to sign with offline keys—or keys sitting off the repository—the policies themselves that confer trust on those packages in the first place (i.e., the root and targets metadata). This ensures that even if an attacker compromises the TUF repository, they cannot tamper with policies without being detected because they do not have access to offline keys, and so they cannot tamper with packages because their signatures would not match the expected policy. Note that this idea generalizes to different policies for different packages in a shared repository, although the exact details are left to future work.
Figure 3: DiVerify uses TUF[43] to distribute configuration and policy securely via a chain of roles. The Root role establishes trusted keys for other roles, the Targets role signs metadata about the policy, and the Snapshot and Timestamp roles ensure integrity and freshness of the repository state.
TUF requires developers to bootstrap the repository with an offline key ceremony, during which they define which keys or identities are allowed to sign which pieces of metadata (notably about DiVerify policies). After this one-time setup, developers need only sign and publish new TUF targets metadata about new DiVerify policies, and machines can be safely trusted to sign and publish only the “news” whether there are new policies at all.
When verifying a package signature, DiVerify first uses TUF to fetch and verify the latest policy. This includes validating the full signature chain, starting from the root metadata down to the targets metadata that points to the exact policy. Only after this verification succeeds does DiVerify apply the now-verified policy to evaluate the package’s signing proofs.
6 Adapting DiVerify to Existing Systems
@@
Figure 4: Two deployment modes for DiVerify. (a) Diverse verification without client integrity attestation. (b) Diverse verification with client integrity attestation. The latter provides stronger security by ensuring that verification results cannot be tampered with by a compromised client.
While DiVerify defines a core trust model for diverse identity verification, tool, and process integrity, the architecture admits multiple realizations with varying security guarantees. DiVerify can be incrementally adapted to fit into existing identity-based signing systems by adjusting trust assumptions, replacing components, or selectively omitting features like attestation or trusted hardware.
This section outlines how DiVerify can be instantiated in different environments, each offering tradeoff between security and deployability. These variants demonstrate how DiVerify’s underlying mechanisms can integrate with real-world existing signing solutions that operate under different constraints.
6.1 Legacy-Compatible
@@
Current solutions typically rely on a single authentication mechanism - an OIDC identity token. In this variant as shown in Figure˜4a, DiVerify operates without a secure daemon. Instead, the client is trusted to honestly retrieve and present authentication claims, and the DiVerify proof becomes a collection of such claims, embedded directly into an existing root of trust (e.g., a signing certificate).
@@
Even in this relaxed model, software maintainers can define DiVerify policies that go beyond a single identity source, and constraint for such identities to sign.
However, this setting sacrifices client-side integrity guarantees. Because there is no trusted daemon or attestation, the system cannot validate client integrity. A compromised client could spoof or replay scopes. While a certificate authority may verify the authenticity of claims against the issuing scope providers, this only holds for secure and verifiable providers that are able to authenticate themselves.
6.2 Trusted Authentication
The Trusted Authentication variant (depicted in  Figure˜4b) is similar to the core DiVerify model but differs in when the daemon’s integrity is verified and who performs the verification.
As in the core model, the daemon retrieves scopes from scope providers. However, instead of proceeding to signing after collecting and validating scopes, the daemon defers signing until an external verifier attests to its integrity. This additional check helps prevent a compromised daemon from producing valid signatures.
A common choice for this external verifier is a Certificate Authority, which can embed the result of the daemon’s attestation into a signing certificate. This approach aligns with existing infrastructures such as those described in [62], where CAs verify the identity of a signer before issuing a certificate. However, the external verifier does not need to be limited to traditional CAs. Alternatives include federated identity systems, policy engines, or custom services that can enforce integrity checks before authorizing a signing operation.
By requiring the daemon’s integrity to be confirmed prior to signing, this variant provides a proactive safeguard against signature generation by a compromised daemon. However, this setting still relies on a trusted CA deployment, which introduces a single point of failure.
7 Security Evaluation
7.1 Security Analysis
@@
Certificate Authority $Circle \CIRCLE -$
We consider three attack scenarios in identity-based signing, derived from our threat model (section˜3.2) and show how our core, legacy-compatible section˜6.1 and trusted authentication section˜6.2 DiVerify variant protects against them.
First, the compromise of the scope provider to issue a fraudulent signer scope; second, the compromise of the client to subvert the signing process and sign malicious artifacts; and third, the compromise of the certificate authority to conceal a malicious client.
Scope Provider Compromise (Types I & II)
If a scope provider were to be compromised (say, by hacking an OIDC provider), they can issue fraudulent scopes to impersonate a legitimate signer.
Under our system assumption, even if an attacker fraudulently obtains one valid identity token (say, by hacking an OIDC provider), they still lack the other required scopes – for example, they cannot fake the TEE-based attestation from the signing client. Because at least one of the multiple checks will fail, the attacker cannot produce a complete DiVerify proof to sign malicious code. Hence, a single compromised provider is not enough to break the chain of trust, thus providing security goal S1 and S2 guarantees
Client Compromise (Type III)
@@
DiVerify’s resilience depends on the operating mode. In Legacy-Compatible mode, the client is trusted to authenticate signers, but without a TEE its behavior cannot be checked. A compromised client can spoof or omit verification steps, leaving this mode vulnerable.
@@
In contrast, Core mode and Trusted Authentication mode delegate verification to a daemon running inside a TEE, while the untrusted client only forwards requests. Any modification of the daemon or its execution environment is detectable through attestation. Core mode validates the signing context at verification time, while Trusted Authentication mode validates it at certificate issuance. In both cases, signatures produced by a compromised client are detectable and rejected, thereby satisfying Security Goal S3.
@@
A remaining risk is payload substitution: e.g., Alice asks to sign package foofoo, but the client forwards $foo^{\prime}$ to the daemon. Because authentication is decoupled from content, Alice has no visibility into what is actually signed. This can be mitigated if (i) the daemon displays the artifact information such as the file name and its cryptographic hash directly to Alice through a trusted channel (e.g., using GPU passthrough), without relying on the compromised client or (ii) a scope provider attests to the client’s integrity using TPM ME [32] or IMA [48].
Certificate Authority Compromise (Type III)
@@
When a CA is used, and is compromised, they can issue signing certificates without validating the client’s attestation, or may deliberately fail to flag invalid attestation. This allows a rogue client signing to appear legitimate to downstream verifiers. As a result, malicious software may be signed and distributed with a valid certificate. DiVerify mitigates this risk to varying degrees depending on the operating mode.
@@
Trusted Authentication Mode assumes CA trust but includes the client’s attestation within the certificate. If a verifier re-validates the embedded attestation independently rather than trusting the certificate blindly, it can detect that the CA failed to perform required checks holding CAs accountable. This attack is not applicable in Core Mode, as the CA is removed from the trust chain entirely. Instead, the verifier directly validates the client’s attestation before accepting a signature, eliminating the risk of CA compromise. Legacy-Compatible Mode offers no protection against this attack. Since attestations are not part of the signing process, there is no mechanism for verifying whether the CA observed valid scopes or verified the client’s integrity before issuing the certificate.
7.2 Mitigating Past Attacks with DiVerify
We illustrate how DiVerify could have mitigated the attacks earlier described in section˜3.2.
  1. 1.

### v3

@@ 5.3. Integrating Scope Providers
DiVerify requires that scope providers must be able to authenticate themselves to the daemon and prevent scope replay. Not all providers natively support these properties. To accommodate a wide range of provider capabilities, DiVerify supports two integration strategies; wrapping and TEE-based integration. We explore different provider categories and their are support in DiVerify.
@@
Verifiable Providers (VP) can cryptographically authenticate themselves and issue scopes that are bound to a specific context, making them verifiable and replay resistant. A representative example is an OIDC token signed by a known Identity Provider such as Google. VPs are natively supported in DiVerify; integrating them requires only defining the appropriate syntax for policy verification.
@@
Static Providers (SP) is not replay resistant, but the authenticity of the provider can still be verified through cryptographic signatures or secure communication channels. SPs are often stateless protocols that provide a fixed value whenever a scope is requested. For example, hardware keys using PIV device attestation provide the same X.509 certificate for each request. This design introduces two risks: (1) without state, an attacker can replay cached scope proofs; and (2) fixed values can be reused across contexts, enabling misuse outside the intended signing event.
To support SPs, DiVerify provides two mechanisms:
@@
  • Zk-based SP: The provider is wrapped in a process that produces a zero-knowledge proof attesting that the value corresponds to a valid device fingerprint. This proof can be verified by both the DiVerify daemon and the verifier, and incorporates a nonce synchronized with other scope providers.
  • nonce-hash SP: In cases where authentication suffices and only event binding is required, the provider output is hashed together with a nonce. The provider returns a tuple (noncenonce, measurementmeasurement, sigsig(hashhash(noncenoncemeasurementmeasurement)))
@@
Opaque Providers (OP) allow nonce injection to prevent replay attacks, but lack a reliable way to verify their identity, making them susceptible to spoofing if used in isolation. For example, this is the case of TPM ME and device fingerprinting writ large.
@@
Untrusted Providers (UT) support neither nonce injection nor provider verification, offering no protection against replay or impersonation attacks and representing the weakest security posture. We combine the strategies for OPs and SPs to support UTs.
@@
6. Policy Semantics and Lifecycle
As discussed in section 5.2, verifiers rely on policies to decide when a signature is trustworthy. They must interpret evidence under consistent trust expectations that can evolve over time. This section describes how DiVerify encodes and securely updates those expectations.
6.1. DiVerify Policy Language
DiVerify policies allow software maintainers to state the trust conditions under which a signature should be accepted. Signing clients produce proofs that contain authenticated scopes and an integrity attestation of the signing tool. A verifier must determine whether this evidence satisfies the expectations expressed in the policy.
These expectations vary across settings. For example, a signer working on a personal computer may authenticate with an identity provider, present a hardware token, and show that signing occurred on a specific trusted device. When signing from a remote environment, the same signer may still prove identity or possession of a hardware key but may no longer be able to prove use of the expected device. DiVerify must support this practical flexibility. Verifiers should be able to express strict requirements when needed, as well as fallback trust paths when some contextual evidence is unavailable. Since DiVerify aggregates identity, contextual, and attestation evidence into a structured proof, verifiers require a policy language that can express such trust requirements in a deterministic and verifiable manner.
The DiVerify policy language is designed to meet these goals. First, it aligns with the structure of the DiVerify proof so that verifiers can evaluate evidence without ambiguity. Second, it is expressive enough to capture realistic signing workflows, including cases where different forms of authentication may be acceptable. Third, it is safe: a policy cannot be satisfied by combining claims from multiple signers or from inconsistent signing contexts.
To achieve this, a policy specifies constraints over three categories of evidence produced during proof construction: authenticated scopes, attestation properties of the signing environment, and a logical rule that describes how these claims must combine to satisfy the maintainer’s trust requirements.
Formally, a policy is a tuple:
$P=\langle Scope,Attest,Rule\rangle$
Scope describes what authentication claims may exist in the signature’s DiVerify proof. It is defined per signer and is structured as a collection of typed claims. For a signer $u$, the scope is modeled as:
$Scope(u)=\{\langle t_{1},v_{1}\rangle,\langle t_{2},v_{2}\rangle,\ldots,\langle t_{k},v_{k}\rangle\}$
where each $t_{i}$ denotes a scope type and $v_{i}$ is the value to be asserted by the corresponding Scope Provider.
Because a signer may possess multiple values for a given scope type (e.g., multiple trusted devices), scopes are modeled as a mapping from scope types to sets of acceptable values. For a signer $u$, $Scope(u)[t]=\{v_{1},\ldots,v_{n}\}$ is the set of values verifiers can trust. At verification time, Let $\Sigma=\{\sigma_{1},\ldots,\sigma_{n}\}$ be the set of different scopes contained in a DiVerify proof. A scope predicate $\textsf{hasScope}(t,x)$ evaluates to true if there exists a scope $\sigma\in\Sigma$ of type $t$ whose value equals $x$.
AttestAttest captures verifier-side trust requirements over the integrity of the signing environment and is modeled as a tuple
$Attest=\langle M,B,Q\rangle$
where $M$ is the expected environment measurement, $B$ is the binding value that cryptographically links the attested environment to the signing key and the authenticated scopes, and $Q$ specifies the verification procedure for TEE quotes. Given an attestation report $q$, the predicate $\textsf{validAttest}(q,\mathit{Attest})$ holds if the quote verifies under $Q$, the included measurement matches $M$, and the binding value contained in $q$ matches $B$ as derived from the signing key and authenticated scopes.
The $Rule$ defines the conditions under which a verifier considers the provided scopes and attestation evidence sufficient. It merges two concerns: (1) value-level checks ensuring a match with expected values and (2) logical composition specifying how multiple signer claims must relate to one another. A semantic constraint is that all predicates within a Rule refer to the same signer. That is, Rule expressions do not compose evidence across different signers; the verifier evaluates whether a subject signer consistently meets all required conditions. We express these as a Boolean expression over scope assertions:
$Rule=$ $\textsf{match}(u,t,x)\mid Rule_{1}\land Rule_{2}\mid Rule_{3}\lor Rule_{4}\mid\lnot Rule_{5}$
which evaluates to true when signer $u$ has a scope of type $t$ whose value set contains $x$. Complex Rules are formed by combining these predicates using the logical operators AND, OR, and NOT. All predicates in the Rule refer to the same signer; a Rule is satisfied only if there exists a single signer $u$ for which the entire Boolean expression evaluates to true.
6.2. Policy Metadata, Update, and Revocation
In addition to the semantic content of the policy which specifies the required identity claims, attestation properties, and Rule evaluation, the system must also represent how policies evolve over time and how verifiers determine whether the policy they are using is the most current and authoritative. To support secure policy evolution, we augment each policy with a lightweight layer of policy metadata. This metadata enables verifiers to authenticate the policy, detect replayed or stale versions, and reason about both semantic updates and revocation events.
A policy instance is therefore modeled as a pair $(Meta,P)$ where P is the semantic policy described in section 6.1 and Meta is a signed metadata record that governs its validity and has the form:
$Meta=\langle\textsf{version},\textsf{epoch},\textsf{issuedAt},\textsf{expiresAt},h_{P},sig_{pol},sig_{root}\rangle$
where $h_{P}$ is the hash of policy, binding the metadata to a specific policy. The version number is monotonically increasing integer that reflects semantic updates to the policy (as shown in Figure 3). Any changes that affect the meaning of the policy, such as updating trusted signer, updating the scope requirement, or altering the Rule structure, results in an increment to the version. Because such updates change the trust configuration, the metadata for a new semantic update must be jointly authorized by the long-term policy key and a higher-privileged root key. This dual signature ($sig_{pol},sig_{root}$) requirement ensures that no single actor can unilaterally introduce or weaken trust conditions.
Figure 3. Policy metadata evolution. Vertical changes correspond to version updates that alter policy semantics, while horizontal changes correspond to epoch updates that only refresh metadata.
Not all policy updates modifies semantics. In practice, a policy issuer may need to refresh operational properties of the metadata, such as extending its validity interval or reissuing it to prevent replay, without changing the meaning of the policy. To support these non-semantic updates, the policy metadata incorporates an epoch: another monotonically increasing integer that associates a particular semantic version. Incrementing the epoch yields a new instance of metadata that binds to the same policy $P$ and has the same versionversion (as shown in Figure 3). Because epoch increments do not modify the policy’s trust semantics, they require only the signature of the current key.
The policy key defines the authority permitted to issue metadata for a given semantic version. For this reason, rotating the policy key itself is a semantic change as it changes that authority. DiVerify therefore treats policy key rotation as a semantic update that must be co-authorised by the root key.
The metadata timestamps $issuedAt$ and $expiresAt$ provide verifiers with freshness guarantees. A verifier accepts a policy if its metadata only if the current time lies within its validity intervals. These timestamps prevent an adversary from indefinitely replaying an older signed metadata record. Verifiers maintain, for each policy identifier, the highest (version, epoch) pair they have previously accepted. Upon receiving new metadata, the verifier checks its signature and validates the hash binding to $P$. It then compares the incoming metadata against its stored maxima. Any metadata record with a lower version or with a lower epoch for the same version is rejected as a rollback attempt. This small amount of persistent state suffices to ensure strong rollback resistance in offline and adversarial settings, while imposing negligible storage overhead. DiVerify requires significantly less state than the manifest and metadata caches routinely maintained by software verification tools, such as APT (Hertzog and Mas, 2014), DNF (Red Hat, [n.d.]), and TUF (gittuf, 4 29) clients. In deployments where such systems already manage trusted metadata, DiVerify’s policy metadata can be integrated or subsumed into their existing mechanisms for distribution and caching (section 7).
To handle signer compromise, policy error, and evolving trust requirements, the metadata layer provides a systematic mechanism for revocation. When the policy issuer intends to remove a previously trusted element from the policy, he publishes a new semantic version that encodes this change. Such modification alters the trust semantics and therefore requires co-authorization by the root key. Revocation may also occur implicitly such that if a new semantic version excludes or tightens conditions that earlier versions permitted, then verifiers interpreting the latest version will reject proofs that depend on the superseded trust assumptions. Because verifiers enforce monotonicity over version numbers, any trust configuration that has been superseded cannot be reinstated by replaying older metadata.
7. DiVerify in Existing Systems
@@
Figure 4. Two integration models for using DiVerify with existing signing systems, in addition to the Core Model ( Figure 2). (a) Legacy-Compatible (no client attestation). (b) Trusted Authentication (with client attestation). The latter protects against compromised clients.
While DiVerify defines a core trust model for identity and client verification, its architecture supports multiple realizations with varying strength of guarantees. This flexibility enables incremental adoption: existing systems can integrate DiVerify using available capabilities and progressively enforce stronger trust requirements as they become ready. In this section, we describe two DiVerify instantiations that illustrate how the same underlying mechanisms integrate with existing signing systems operating under different operational constraints.
7.1. Legacy-Compatible
@@
Current solutions typically rely on a single authentication mechanism: an OIDC identity token. In this variant as shown in Figure 4a, DiVerify operates without a secure daemon. Instead, the client is trusted to honestly retrieve and present authentication claims, and the DiVerify proof becomes a collection of such claims, embedded directly into an existing root of trust (e.g., a signing certificate).
@@
Even in this relaxed model, software maintainers can define DiVerify policies that go beyond a single identity source, and constraint for such identities to sign. However, this setting sacrifices client-side integrity guarantees. Because there is no trusted daemon or attestation, the system cannot validate client integrity. A compromised client could spoof or replay scopes. While a certificate authority may verify the authenticity of claims against the issuing scope providers, this only holds for secure and verifiable providers that are able to authenticate themselves.
7.2. Trusted Authentication
The Trusted Authentication variant (depicted in  Figure 4b) is similar to the core DiVerify model but differs in when the daemon’s integrity is verified and who performs the verification.
As in the core model, the daemon retrieves scopes from scope providers. However, instead of proceeding to signing after collecting and validating scopes, the daemon defers signing until an external verifier confirms its integrity. This additional check helps prevent a compromised daemon from producing valid signatures.
A common choice for this external verifier in existing systems is a Certificate Authority, which can embed the daemon’s attestation into a signing certificate. This approach aligns with existing infrastructures such as those described in (Newman et al., 2022), where CAs verify the identity of a signer before issuing a certificate. However, the external verifier does not need to be limited to traditional CAs. Alternatives include federated identity systems, policy engines, or custom services that can enforce integrity checks before authorizing a signing operation.
By requiring the daemon’s integrity to be confirmed prior to signing, this variant provides a proactive safeguard against signature generation by a compromised daemon. However, this setting still relies on a trusted CA, which introduces a single point of failure.
8. Security Analysis
@@
Certificate Authority $Circle \CIRCLE$
We consider three attack scenarios (Table 1) in identity-based signing, derived from our threat model (section 3.2) and show how our core, legacy-compatible (section 7.1) and trusted authentication (section 7.2) DiVerify variants protect against them.
Types I & II (Scope Provider Compromise)
If a scope provider were to be compromised, an attacker can issue fraudulent scopes to impersonate a legitimate signer.
Under this assumption, even if an attacker fraudulently obtains one valid identity token, they still lack the the ability to provide further scopes (e.g., they cannot fake the TEE-based attestation from the signing client). Because at least one of the multiple checks will fail, the attacker cannot produce a complete DiVerify proof to sign malicious code. Hence, a single compromised provider is not enough to break the chain of trust, thus providing security goal S1 and S2 guarantees.
Type III (Client Compromise)
@@
DiVerify’s resilience depends on the operating mode. In Legacy-Compatible, the client is trusted to authenticate signers, but without a TEE its behavior cannot be checked. A compromised client can spoof or omit verification steps, leaving this mode vulnerable.
@@
In contrast, Core mode and Trusted Authentication mode delegate verification to a daemon running inside a TEE, while the untrusted client only forwards scope requests. Any modification of the daemon or its execution environment is detectable through attestation. Core mode validates the signing context at verification time, while Trusted Authentication mode validates it at certificate issuance. In both cases, signatures produced by a compromised client are detectable and rejected, thereby satisfying Security Goal S3.
@@
A remaining risk is payload substitution: e.g., Alice attempts to sign package foo, but the client forwards $\textit{foo}^{\prime}$ to the daemon. Because authentication is decoupled from content, Alice has no visibility into what is actually signed. This is mitigated if (i) the daemon displays the artifact information such as the file name and its cryptographic hash directly to Alice through a trusted channel (e.g., using GPU passthrough), without relying on the compromised client or (ii) a scope provider attests to the client’s integrity using TPM ME (Intel Corporation, 2025) or IMA (Linux Integrity Subsystem, 4 21).
Certificate Authority Compromise
@@
When a CA is used (as in Legacy-Compatible and Trusted Authentication modes), and is compromised, they can issue signing certificates without validating the client’s attestation, or may deliberately fail to flag invalid attestation. This allows a rogue client signing to appear legitimate to downstream verifiers. As a result, malicious software may be signed and distributed with a valid certificate. DiVerify mitigates this risk to varying degrees depending on the operating mode.
@@
Trusted Authentication mode assumes trusted CA but includes the client’s attestation within the certificate. If a verifier re-validates the embedded attestation independently rather than trusting the certificate blindly, it can detect that the CA failed to perform required checks, holding CAs accountable. This attack is not applicable in Core mode, as the CA is removed from the trust chain entirely. Instead, the verifier directly validates the client’s attestation before accepting a signature, eliminating the risk of CA compromise. Legacy-Compatible mode offers no protection against this attack. Since attestations are not part of the signing process, there is no mechanism for verifying whether the CA observed valid scopes or verified the client’s integrity before issuing the certificate.
8.1. Mitigating Past Attacks with DiVerify
We illustrate how DiVerify mitigates the attacks described in section 3.2.
  1. (1)